### v1

@@ 3.6 Co-added spectra
For a more robust spectral classification, we produced high S/N spectra by stacking the nine available epochs for each target. To achieve this, the cross-correlation occurred in two steps. First, the spectra were shifted to the rest frame by cross-correlating them with a spectral model and then co-added, which enabled us to generate a first co-added spectrum. In this step, the entire spectral range was considered, which is dominated by the Balmer lines H and H. As a second step, we cross-correlated all the epochs against the co-added spectrum formed in the previous step, which is calibrated to the rest frame. This time, however, we used a narrow spectral window around He i (namely 4460–4485 Å), which is present in essentially all objects555For AF supergiants, this range is rich in other spectral lines such as Mg ii , making the cross-correlation in this range equally possible.. This allows for a more accurate radial-velocity (RV) measurement that is tuned to the spectral morphology of the individual star. For the first step, we used a pre-computed model from the TLUSTY O-star model atmosphere grid (Hubeny & Lanz, 1995; Lanz & Hubeny, 2003) for 666This is the closest available metallicity to that of the SMC. In any case, the impact of the metallicity on RV measurements is negligible, an effective temperature of kK, and a surface gravity of []. The exact choice of the model can impact the absolute RV calibration, but has no impact on the process of spectral classification which is the focus of this paper.
We note that for double-lined spectroscopic binaries (SB2), the process of co-adding the data will smear the spectral features of the two components and result in an average spectrum, which may well contaminate the classification. Clear cases of SB2s (40/929 in total) were identified via visual inspection of the available epochs (see also Sect. 4). A montage of the co-added spectra of selected stars, ordered by descending spectral type (Sect. 4), is shown in Figs. 4 – 7.
@@ 4 Spectral classification
We have established SMC reference OB stars from comparison with Galactic O star templates from Sota et al. (2011) and Maíz Apellániz et al. (2016) or Galactic B star templates from Negueruela et al. (2024). These stars are drawn from BLOeM datasets, supplemented by archival VLT/FLAMES (Evans et al., 2006; Dufton et al., 2019) or VLT/X-Shooter (Vink et al., 2023) datasets. Reference stars are ideally sharp-lined, permitting rotational broadening to be applied for comparison with fast rotators. We assign e, e+ and pe for stars exhibiting (i) H i, (ii) H i & Fe ii, and (iii) H i & He i emission, respectively. The qualifier ”:” deems the classification uncertain. We note that the BLOeM spectral range includes most diagnostic lines, but misses a few lines which can help refine the classification such as N iii , He ii , H, and H.
@@
Spectral types of O to B0.7 stars are determined following the Galactic O star scheme of Sota et al. (2011) and Maíz Apellániz et al. (2016) which utilises the ratio of He ii 4542 to He i 4471 or 4388 for late subtypes (O8.5+). Luminosity classes are usually assigned courtesy of He ii 4686. Since this line is not available for the BLOeM dataset, luminosity classes for O4–7 stars are estimated from the detection of N iv 4058 and Si iv 4088-4116 emission lines. For late O and early B subtypes we use a combination of H and H line profile morphologies supplemented by the ratio of Si iv 4088, 4116 to He i 4121, 4144 lines, following Walborn & Fitzpatrick (1990). It is well known that incorrect luminosity classes would be assigned for silicon-poor SMC stars based on Galactic templates (Walborn, 1983). For example, Walborn et al. (2014) showed that using Si-He criteria results in different luminosity classes than He-line criteria alone. Hence, SMC reference stars are essential for this approach. Nomenclature linked to the region of He ii 4686 such as f and (f), which reflect the presence of emission in the line, is not possible, but we are able to flag potential rapidly rotating O stars via (n), n, nn etc. on the basis of the FWHM of He ii 4542 (early and mid O-types) and He i 4471 (late O-types). For a few O-type stars, spectral types (mostly luminosity classes) were adjusted based on previous literature, given the lack of diagnostics such as He ii and H in the BLOeM dataset. These cases are documented in the comments in Table LABEL:tab:SpT.
@@
We follow the SMC B-type scheme of Lennon (1997) in assigning spectral types of B1–9 stars from Si iv 4088–4116, Si iii 4553, Si ii 4128–32, plus the ratio of He i 4471 to Mg ii 4481 at late subtypes. Luminosity classes are obtained from line morphologies of H and H with respect to reference stars. Unusual B-type systems with Balmer emission lines and forbidden [Fe ii] lines are classified as supergiant B[e] (sgB[e]) following Lamers et al. (1998); Kraus (2019).
@@
For A and F supergiants we are unable to follow the Ca ii H+K criteria of Evans & Howarth (2003) since our dataset excludes the Fraunhofer K line, so again we have selected SMC reference A0–5 supergiants from archival VLT/FLAMES spectroscopy, with luminosity classes assigned from the equivalent width of H from comparison with Millward & Walker (1985). Spectral types assigned to late A and F stars are relatively coarse, and are primarily based on the strength of the CH G-band in BLOeM datasets spectrally degraded to the resolution of SMC AF supergiants presented by Evans & Howarth (2003).
@@ 5.1 Census
A histogram of the spectral type distribution across the range O4 to F5, subdivided into dwarfs/giants and supergiants, is shown in Fig. 8. The BLOeM sample includes the complete spectral range from the earliest stars in the SMC (excluding Wolf-Rayet stars) to yellow supergiants. Overall, the sample includes 159 O-type stars, 324 early B-type dwarfs and giants (B0 – B3 V–III), 309 early B-type supergiants (B0 – B3 II–I), and 137 late-type supergiants (B4 and later). Of these, 75 objects are classified as Oe/Be stars, portraying characteristic emission in their Balmer lines: 20 Oe stars (including four uncertain cases, marked as ”e?”), and 55 Be stars (Sect. 5.2).
@@
The number of dwarfs/giants (generally interpreted as main-sequence objects) steeply decreases around a spectral type of B2, which roughly corresponds to (Harmanec, 1988). Among the B1-B2 dwarfs/giants, the vast majority are giants, bright enough to have been included in our survey. These are likely evolved main sequence stars. In the O-star regime, we are sensitive to the full extent of the main sequence down to the youngest stars, assuming those are not dust-enshrouded.
5.2 Oe/Be fraction: complementary Gaia low-resolution spectra
@@
As noted in Sect. 5.1, 75 targets have been identified as Oe/Be stars in our sample. However, the BLOeM dataset doesn’t cover the H line, which is the most sensitive line for circumstellar material in the visual range. Due to this, it is likely that we have missed Oe/Be stars in the sample which exhibit H emission but not H, which is covered by BLOeM. To mark additional Oe/Be candidates, we make use of Gaia XP low-resolution flux-calibrated spectra (De Angeli et al., 2023), which exist for all sources except BLOeM 4-041, 6-044, 7-012 and 8-033. While these data are of very low resolution (), and some are affected by problematic wiggles in the data, they may be used to detect the presence of strong emission, or candidate OBe stars in our case. To be pragmatic, we adopt a straightforward approach of measuring the H equivalent width (EW) in the XP ‘sampled’ data, using broad windows on both sides of the line to define a local continuum. While more complex approaches are possible (e.g., Weiler et al., 2023), our approach should be adequate for detecting strong H emission. Candidate OBe stars are then flagged as 1- outliers in the distribution of equivalent width as a function of magnitude. As a check on this approach, a visual inspection of all spectra was also performed to define a second category of ‘by eye’ OBe candidates. Both samples are illustrated in Fig. 9. Both methods agree for most targets, although given the contamination induced by the spectral wiggles on the EWs, we favour the by-eye classification of H emitters.
@@
Overall, in addition to the 75 Oe/Be stars classified using the BLOeM dataset, we identify 28 additional H emitters. However, the majority of those are supergiants, which are not Oe/Be stars per definition, or stars with identified nebular contamination. Only seven targets are potential new Be stars. Adding those, the total Oe/Be fraction among the non-supergiant OB stars is 11%. This fraction is much lower than the fraction of 30% that has been found in previous work (Bonanos et al., 2010; Schootemeijer et al., 2022). The difference can be explained by the diagonal CMD cut that we use to select our sample, which favours blue stars and hence disfavours OBe stars, which are about 0.3 magnitudes redder than OB stars (Schootemeijer et al. 2022, their figure A.5).
The bias mentioned above likely does not impact the Oe fraction derived here, which is 20/159 = 13%, comparable to the 10% fraction reported by Bonanos et al. (2010). However, the Be fraction among early B-type stars in the sample (removing supergiants) is 55/560 = 10%, which is likely strongly underestimated due to the exclusion of reddened stars.
@@ 5.3 Notable targets
We highlight a few unique objects in our sample. BLOeM 2-116, BLOeM 3-012, and BLOeM 4-055 are classified as sgB[e] stars. BLOeM 2-116 (alias LHA 115-S 18) was subject of several studies (e.g., Shore et al., 1987; Clark et al., 2013; Maravelias et al., 2014), as was BLOeM 3-012 (alias RMC 4; Zickgraf et al. 1996; Graus et al. 2012; Pasquali et al. 2000; Wu et al. 2020). Such objects are thought to represent a brief evolutionary phase of massive stars. Their spectra resemble those of luminous blue variables (LBVs), and their origin is debated in the literature (e.g. Podsiadlowski et al., 2006; Clark et al., 2013). The spectra of all these objects indicate variability. Whether this variability stems from binary motion is not yet clear, though BLOeM 3-012 has previously been reported to be a binary (Zickgraf et al., 1996) and seems well explained as the product of a binary merger in a triple system (Pasquali et al., 2000; Podsiadlowski et al., 2006; Wu et al., 2020).
BLOeM 3-031 and BLOeM 5-071 show spectra which resemble those of Be + bloated stripped-star binaries such as LB-1 (Irrgang et al., 2020; Abdul-Masih et al., 2020; Shenar et al., 2020; El-Badry & Quataert, 2021) and HR~6819 (Bodensteiner et al., 2020; Frost et al., 2022). Such objects feature strong Balmer emission characteristic of Be stars, in combination with narrow and RV variable absorption features which stem from a putative bloated stripped-star companion.
BLOeM 2-104 and BLOeM 4-039 were classified as Of?p stars by Evans et al. (2004), which refers to the presence of emission lines associated with N iii and C iii in the range 4630–4660Å (Walborn, 1972). We cannot independently verify this since the BLOeM data lack this range, thoug we do confirm the presence of emission in the Balmer lines characteristic of such stars, and so we adopt this classification. All known Galactic Of?p stars possess strong, global magnetic fields (Grunhut et al., 2017). It is therefore plausible that BLOeM 2-104 and BLOeM 4-039 are magnetic as well, although spectropolarimetric data are needed to verify this.
Nine targets are identified as significant X-ray emitters from a cross-match with various X-ray catalogues described in Appendix B, one of which may be a spurious X-ray detection (BLOeM 6-116). Four of those have been classified as high-mass X-ray binaries (HMXBs) in the past: BLOeM 2-055, 2-82, BLOeM 2-116, BLOeM 4-026 and BLOeM 4-113. BLOeM 3-042 may be a colliding-wind binary (CWB) given its classification (O6 III + O7.5). The origin of X-rays in BLOeM 8-029 (B1 IV:) and BLOeM 1-102 (O6 III(n)) is not yet clear. More details regarding the cross-match process and X-ray luminosities of the objects can be found in Appendix B.
Finally, cross-matching with the OGLE catalogue of photometrically variable stars in the SMC (Pawlak et al., 2016), 74/929 (i.e., 8%) of the BLOeM targets are found to be eclipsing binaries, while eight are classified as ellipsoidal variables (Sect. B.4).
@@ 5.4 Hertzsprung–Russell Diagram
Bolometric luminosities are estimated from Ks band point spread function (PSF) photometry fitting from VISTA/VMC (Cioni et al., 2011), a distance modulus of 18.95 mag for the SMC (Hilditch et al., 2005; Graczyk et al., 2020), intrinsic V-Ks colours from Martins & Plez (2006), Cox (2000) and Koornneef (1983), plus V-band bolometric corrections drawn from Bouret et al. (2013) and Bouret et al. (2021) for O stars, Lanz & Hubeny (2007) for B stars with effective temperatures 15kK, Kudritzki et al. (2008) for late B and A supergiants (8kK 12kK) and Cox (2000) otherwise. The use of Ks band photometry avoids the need to correct for interstellar extinction, which is significantly smaller at the Ks band than at the band. Given this and the typically negligible extinction towards SMC sightlines (see, e.g., Schootemeijer et al., 2021), we assume here; this may affect the luminosities of individual objects, but does not affect our general conclusions.
@@
We present the resulting Hertzsprung–Russell diagram (HRD) of the BLOeM sample in Fig. 10, together with evolutionary tracks and isochrones for SMC metallicity stars from the extended grid of Schootemeijer et al. (2019). As described in Appendix B of Hastings et al. (2021), these models have efficient semiconvection () and mass-dependent overshooting ( linearly increases from 0.1 at to 0.3 at 20 , and remains constant at 0.3 for higher masses). Apart from and , this grid has the same physics assumptions as Brott et al. (2011). We adopt the dwarf temperature calibration for sub-giants, and the giant temperature calibration for bright giants. The BLOeM sample includes dwarfs and sub-giants in the mass range 10–70 , plus a few giants and supergiants down to 7–8 . The three sgB[e] objects (BLOeM 2-116, BLOeM 3-012, and BLOeM 4-055) are excluded from the HRD, since standard calibrations cannot be used on them.
@@
The errors on and can only be roughly estimated. The spectral types can be estimated to within one spectral bin roughly, corresponding to ; this is larger than typical calibration errors. However, the impact of binary contamination is not possible to quantify without further analysis; we set the error in Fig. 10 to 0.03 dex. Similarly, errors on bolometric correction can be estimated from neighbouring spectral-type bins. Depending on the spectral type, they typically range between , corresponding roughly to dex. Added to this are potential sources contaminating the Ks magnitudes (e.g., infrared excess), binary contamination, and reddening. To remain conservative, we adopt an uncertainty of dex in Fig. 10.
@@
The HRD reveals the span of parameters covered by our sample in terms of initial mass and age. Qualitatively, it is similar to that presented by Humphreys & McElroy (1984), though the BLOeM sample goes substantially deeper in magnitude (mag). This HRD suggests that the sample probes the range , though only a minority of targets exceed initial masses . In terms of ages, a visual comparison with isochrones implies that the BLOeM sample probes stars as young as Myr, extending to ages of the order of 20 Myr or more. The lack of very massive stars in the SMC was noted in the past (e.g., Ramachandran et al., 2019; Schootemeijer et al., 2021), and could be related to a peak of star formation Myr ago (Antoniou et al., 2010; Rubele et al., 2015; Schootemeijer et al., 2021). The Humphreys–Davidson (H-D) limit (Humphreys & Davidson, 1979; Humphreys & McElroy, 1984), which marks the absence of bright stars in the upper-right part of the HRD, is clearly seen above , in agreement with recent evaluations of the H–D limit in the SMC (Ramachandran et al., 2019; Davies et al., 2018; Gilkis et al., 2021; Sabhahit et al., 2021).
@@
We note that these estimates are subject to uncertainty given the uncertainties on , and could also change depending on the set of evolutionary tracks being used (e.g., Georgy et al., 2013; Choi et al., 2016; Marigo et al., 2017; Keszthelyi et al., 2022). Moreover, an unknown fraction of the sample stars will be affected either by past binary interactions or bright companions, while we only make use of single-star tracks in this first characterisation effort. A full analysis of the mass and age distribution will require treatment of multiplicity (e.g., binary identification, orbital analysis, spectral disentangling), which will be the subject of subsequent papers.

### v2

@@ 3.6 Co-added spectra
For a more robust spectral classification, we produced high-S/N spectra by stacking the nine available epochs for each target. To achieve this, the cross-correlation occurred in two steps. First, the spectra were shifted to the rest frame by cross-correlating them with a spectral model and then co-added, which enabled us to generate a first co-added spectrum. In this step, the entire spectral range was considered, which is dominated by the Balmer lines H and H. As a second step, we cross-correlated all the epochs against the co-added spectrum formed in the previous step, which is calibrated to the rest frame. However, this time we used a narrow spectral window around He i (namely 4460–4485 Å), which is present in almost all objects555For AF supergiants, this range is rich in other spectral lines, such as Mg ii , making the cross-correlation in this range equally possible.. This allows a more accurate radial-velocity (RV) measurement that is tuned to the spectral morphology of the individual star. For the first step, we used a precomputed model from the TLUSTY O-star model atmosphere grid (Hubeny & Lanz, 1995; Lanz & Hubeny, 2003) for 666This is the closest available metallicity to that of the SMC. In any case, the impact of the metallicity on RV measurements is negligible, an effective temperature of kK, and a surface gravity of []. The exact choice of the model can impact the absolute RV calibration, but has no impact on the process of spectral classification, which is the focus of this paper.
We note that for double-lined spectroscopic binaries (SB2), the process of co-adding the data will smear the spectral features of the two components and result in an average spectrum, which may well contaminate the classification. Clear cases of SB2s (40/929 in total) were identified via visual inspection of the available epochs (see also Sect. 4). A montage of the co-added spectra of selected stars ordered by descending spectral type (Sect. 4) is shown in Figs. 4 – 7.
@@ 4 Spectral classification
We established SMC reference OB stars from comparison with Galactic O star templates from Sota et al. (2011) and Maíz Apellániz et al. (2016) or Galactic B star templates from Negueruela et al. (2024). These stars are drawn from BLOeM datasets, supplemented by archival VLT/FLAMES (Evans et al., 2006; Dufton et al., 2019) or VLT/X-Shooter (Vink et al., 2023) datasets. Reference stars are ideally sharp-lined, permitting rotational broadening to be applied for comparison with fast rotators. We assign e, e+, and pe for stars exhibiting (i) H i, (ii) H i and Fe ii, and (iii) H i and He i emission, respectively. Objects whose spectra are contaminated by nebular emission are designated ‘neb’. In some cases, it was difficult to discern between intrinsic source emission and nebular contamination. To distinguish between these cases, we made use of images acquired with the wide-field Digitized Sky Survey (DSS) 2-red to identify objects that are embedded within nebulous regions. More details are given in Appendix B. The qualifier ‘:’ deems the classification uncertain. We note that the BLOeM spectral range includes most diagnostic lines, but misses a few lines that can help refine the classification, such as N iii , He ii , H, and H.
@@
Spectral types of O stars are determined following the Galactic O star scheme of Sota et al. (2011) and Maíz Apellániz et al. (2016), which utilises the ratio of He ii 4542 to He i 4471 or 4388 for late subtypes (O8.5+). Luminosity classes are usually assigned courtesy of He ii 4686. As this line is not available for the BLOeM dataset, luminosity classes for O4–7 stars are estimated from the detection of N iv 4058 and Si iv 4088-4116 emission lines. For late O subtypes, we use a combination of H and H line profile morphologies supplemented by the ratio of Si iv 4088, 4116 to He i 4121, 4144 lines, following Walborn & Fitzpatrick (1990). It is well known that incorrect luminosity classes would be assigned for silicon-poor SMC stars based on Galactic templates (Walborn, 1983). For example, Walborn et al. (2014) showed that using Si-He criteria results in luminosity classes that are different from those resulting from the use of He-line criteria alone. Hence, SMC reference stars are essential for this approach. Nomenclature linked to the region of He ii 4686 —such as f and (f), which reflect the presence of emission in the line— is not possible, but we are able to flag potential rapidly rotating O stars via (n), n, nn, and so on, on the basis of the FWHM of He ii 4542 (early and mid O-types) and He i 4471 (late O-types). For a few O-type stars, spectral types (mostly luminosity classes) were adjusted based on previous literature, given the lack of diagnostics such as He ii and H in the BLOeM dataset. These cases are documented in the comments in Table LABEL:tab:SpT.
@@
We follow the SMC B-type scheme of Lennon (1997), supplemented by early B templates from Negueruela et al. (2024), in assigning spectral types from Si iv 4088, Si iii 4553, and Si ii 4128–32, plus the ratio of He i 4471 to Mg ii 4481 at late subtypes. Secondary criteria for B0–0.7 stars (Sota et al., 2011), involving the ratio of He i 4388 to He ii 4541 or He i 4144 to He ii 4200, are used for stars with extremely weak Si diagnostics. Luminosity classes are obtained from line morphologies of H and H with respect to reference stars. Unusual B-type systems with Balmer emission lines and forbidden [Fe ii] lines are classified as supergiant B[e] (sgB[e]) following Lamers et al. (1998) and Kraus (2019).
@@
For A and F supergiants, we are unable to follow the Ca ii H+K criteria of Evans & Howarth (2003), because our dataset excludes the Fraunhofer K line, and so again we selected SMC reference A0–5 supergiants from archival VLT/FLAMES spectroscopy, with luminosity classes assigned from the equivalent width of H from comparison with Millward & Walker (1985). Spectral types assigned to late A and F stars are relatively coarse, and are primarily based on the strength of the CH G-band in BLOeM datasets spectrally degraded to the resolution of SMC AF supergiants presented by Evans & Howarth (2003). All F stars have luminosity class II–I owing to the strength of their metallic features (Gray & Corbally, 2009), even at the low metallicity of the SMC.
@@ 5.1 Census
A histogram of the spectral-type distribution across the range O4 to F5, subdivided into dwarfs and giants in one group and supergiants in the other, is shown in Fig. 8. The BLOeM sample includes the complete spectral range from the earliest stars in the SMC (excluding Wolf-Rayet stars) to yellow supergiants. Overall, the sample includes 159 O-type stars, 331 early B-type dwarfs and giants (B0 – B3 V–III), 303 early B-type supergiants (B0 – B3 II–I), and 136 late-type supergiants (B4 and later). Of these, 82 objects are classified as OBe stars, portraying characteristic emission in their Balmer lines: 20 Oe stars (including four uncertain cases, marked ‘e?’), and 62 are Be stars (Sect. 5.2).
@@
The number of dwarfs and giants (generally interpreted as main sequence objects) steeply decreases around a spectral type of B2, which roughly corresponds to (Harmanec, 1988). Among the B1-B2 dwarfs and giants, the vast majority are giants, bright enough to have been included in our survey. These are likely evolved main sequence stars. In the O-star regime, we are sensitive to the full extent of the main sequence down to the youngest stars, assuming those are not dust-enshrouded.
5.2 OBe fraction: Complementary Gaia low-resolution spectra
@@
As noted in Sect. 5.1, 82 targets have been identified as OBe stars in our sample. However, the BLOeM dataset does not cover the H line, which is the most sensitive line for circumstellar material in the visual range. For this reason, it is likely that we have missed OBe stars in the sample. To mark additional OBe candidates, we make use of Gaia XP low-resolution flux-calibrated spectra (De Angeli et al., 2023), which exist for all sources except BLOeM 4-041, 6-044, 7-012, and 8-033. While these data are of very low resolution (), and some are affected by problematic wiggles in the data, they may be used to detect the presence of strong emission, or candidate OBe stars in our case. To be pragmatic, we adopt a straightforward approach of measuring the H equivalent width (EW) in the XP ‘sampled’ data, using broad windows on both sides of the line to define a local continuum. While more complex approaches are possible (e.g. Weiler et al., 2023), our method should be adequate for detecting strong H emission. Candidate OBe stars are then flagged as 1 outliers in the distribution of EW as a function of magnitude. As verification of this approach, a visual inspection of all spectra was also performed to define a second category of ‘by-eye’ OBe candidates. Both samples are illustrated in Fig. 9. The two methods agree for most targets, although given the contamination induced by the spectral wiggles on the EWs, we favour the by-eye classification of H emitters.
@@
Overall, in addition to the 82OBe stars classified using the BLOeM dataset, we identify 16 additional H emitters. However, half of these are supergiants according to our classification from the BLOeM dataset, which, by definition, are not OBe stars. Only eight targets are potential Be stars that have not been classified as such: BLOeM 4-046, 4-096, 5-096, 5-107, 7-008, 7-094, 7-099, and 8-078. Adding those, the total number of OBe stars in our sample is 90, amounting to a fraction among the non-supergiant OB stars of 11%. This fraction is much lower than the fraction of 30% that has been found in previous work (Bonanos et al., 2010; Schootemeijer et al., 2022). The difference can be explained by the diagonal CMD cut that we use to select our sample, which favours blue stars and hence disfavours OBe stars, which are about 0.3 magnitudes redder than OB stars (Schootemeijer et al. 2022, their figure A.5). The bias mentioned above likely does not impact the Oe fraction derived here, which is 20/159 = 13%, comparable to the 10% fraction reported by Bonanos et al. (2010).
@@ 5.3 Notable targets
We highlight a few unique objects in our sample. BLOeM 2-116, BLOeM 3-012, and BLOeM 4-055 are classified as sgB[e] stars. BLOeM 2-116 (alias LHA 115-S 18) has been the subject of several studies (e.g. Shore et al., 1987; Clark et al., 2013; Maravelias et al., 2014), as has BLOeM 3-012 (alias RMC 4; Zickgraf et al. 1996; Graus et al. 2012; Pasquali et al. 2000; Wu et al. 2020). Such objects are thought to represent a brief evolutionary phase of massive stars. Their spectra resemble those of luminous blue variables (LBVs), and their origin is debated in the literature (e.g. Podsiadlowski et al., 2006; Clark et al., 2013). The spectra of all these objects indicate variability. Whether or not this variability stems from binary motion is not yet clear, though BLOeM 3-012 has previously been reported to be a binary (Zickgraf et al., 1996) and seems well explained as the product of a binary merger in a triple system (Pasquali et al., 2000; Podsiadlowski et al., 2006; Wu et al., 2020).
BLOeM 3-031 and BLOeM 5-071 show spectra that resemble those of Be + bloated stripped-star binaries such as LB-1 (Irrgang et al., 2020; Abdul-Masih et al., 2020; Shenar et al., 2020; El-Badry & Quataert, 2021) and HR~6819 (Bodensteiner et al., 2020; Frost et al., 2022). Such objects feature strong Balmer emission characteristic of Be stars, in combination with narrow and RV variable absorption features that stem from a putative bloated stripped-star companion.
BLOeM 2-104 and BLOeM 4-039 were classified as Of?p stars by Evans et al. (2004), which refers to the presence of emission lines associated with N iii and C iii in the range 4630–4660Å (Walborn, 1972). We cannot independently verify this, because the BLOeM data lack this range, though we do confirm the presence of emission in the Balmer lines characteristic of such stars, and so we adopt this classification. All known Galactic Of?p stars possess strong global magnetic fields (Grunhut et al., 2017). It is therefore plausible that BLOeM 2-104 and BLOeM 4-039 are magnetic as well, although spectropolarimetric data are needed to verify this.
Nine targets are identified as significant X-ray emitters from a cross-match with various X-ray catalogues described in Appendix C, one of which may be a spurious X-ray detection (BLOeM 6-116). Four of those have been classified as high-mass X-ray binaries (HMXBs) in the past: BLOeM 2-055, 2-82, BLOeM 2-116, BLOeM 4-026 and BLOeM 4-113. BLOeM 3-042 may be a colliding-wind binary (CWB) given its classification (O6 III + O7.5). The origin of X-rays in BLOeM 8-029 (B1 IV:) and BLOeM 1-102 (O6 III(n)) is not yet clear. Further details regarding the cross-match process and X-ray luminosities of the objects can be found in Appendix C.
Finally, cross-matching with the OGLE catalogue of photometrically variable stars in the SMC (Pawlak et al., 2016), 74 (i.e. 8%) of the 929 BLOeM targets are found to be eclipsing binaries, while 8 are classified as ellipsoidal variables (Sect. C.4).
@@ 5.4 Hertzsprung–Russell diagram
Bolometric luminosities are estimated from Ks band point spread function (PSF) photometry fitting from VISTA/VMC (Cioni et al., 2011), a distance modulus of 18.95 mag for the SMC (Hilditch et al., 2005; Graczyk et al., 2020), intrinsic V-Ks colours from Martins & Plez (2006), Cox (2000), and Koornneef (1983), plus V-band bolometric corrections drawn from Bouret et al. (2013) and Bouret et al. (2021) for O stars, Lanz & Hubeny (2007) for B stars with effective temperatures 15kK, Kudritzki et al. (2008) for late B and A supergiants (8kK 12kK), and Cox (2000) otherwise. The use of Ks band photometry avoids the need to correct for interstellar extinction, which is significantly smaller at the Ks band than at the band. Given this and the typically negligible extinction towards SMC sightlines (see e.g. Schootemeijer et al., 2021), we assume here; this may affect the luminosities of individual objects, but does not affect our general conclusions. For the intermediate luminosity classes IV and II, we use the dwarf and supergiant relations from Table 1, respectively.
@@
We present the resulting Hertzsprung–Russell diagram (HRD) of the BLOeM sample in Fig. 10, together with evolutionary tracks and isochrones for SMC metallicity stars from the extended grid of Schootemeijer et al. (2019). As described in Appendix B of Hastings et al. (2021), these models have efficient semi-convection () and mass-dependent overshooting ( linearly increases from 0.1 at to 0.3 at 20 , and remains constant at 0.3 for higher masses). Apart from and , this grid has the same physics assumptions as Brott et al. (2011). We adopt the dwarf temperature calibration for subgiants, and the giant temperature calibration for bright giants. The BLOeM sample includes dwarfs and subgiants in the mass range 10–70 , plus a few giants and supergiants down to 7–8 . The three sgB[e] objects (BLOeM 2-116, BLOeM 3-012, and BLOeM 4-055) are excluded from the HRD, as standard calibrations cannot be used on them.
@@
The errors on and can only be roughly estimated. The spectral types can be roughly estimated to within one spectral bin, corresponding to ; this is larger than typical calibration errors. However, the impact of binary contamination is not possible to quantify without further analysis; we set the error in Fig. 10 to 0.03 dex. Similarly, errors on bolometric correction can be estimated from neighbouring spectral-type bins. Depending on the spectral type, they typically range between , corresponding roughly to dex. Added to this are potential sources contaminating the Ks magnitudes (e.g. infrared excess), binary contamination, and reddening. To remain conservative, we adopt an uncertainty of dex in Fig. 10.
@@
The HRD reveals the span of parameters covered by our sample in terms of initial mass and age. Qualitatively, it is similar to that presented by Humphreys & McElroy (1984), though the BLOeM sample goes substantially deeper in magnitude (mag). This HRD suggests that the sample probes the range , though only a minority of targets exceed initial masses of . In terms of age, a visual comparison with isochrones implies that the BLOeM sample probes stars as young as Myr, extending to ages of the order of 20 Myr or more. The lack of very massive stars in the SMC has been noted in the past (e.g. Ramachandran et al., 2019; Schootemeijer et al., 2021), and could be related to a peak of star formation Myr ago (Antoniou et al., 2010; Rubele et al., 2015; Schootemeijer et al., 2021). The Humphreys–Davidson (H-D) limit (Humphreys & Davidson, 1979; Humphreys & McElroy, 1984), which marks the absence of bright stars in the upper-right part of the HRD, is clearly seen above , in agreement with recent evaluations of the H–D limit in the SMC (Ramachandran et al., 2019; Davies et al., 2018; Gilkis et al., 2021; Sabhahit et al., 2021). In Appendix D, we show similar HRDs for each of the eight SMC fields, which provide an overview of the stellar content in each field.
@@
We note that these estimates are subject to uncertainty given the uncertainties on , and could also change depending on the set of evolutionary tracks being used (e.g. Georgy et al., 2013; Choi et al., 2016; Marigo et al., 2017; Keszthelyi et al., 2022). Moreover, an unknown fraction of the sample stars will be affected either by past binary interactions or bright companions, while we only make use of single-star tracks in this first characterisation effort. A full analysis of the mass and age distribution will require treatment of multiplicity (e.g. binary identification, orbital analysis, spectral disentangling), which will be the subject of subsequent papers.